# copyright: none

### v2

@@ Abstract.
[…]
[[[[Experiment, Analysis & Benchmark]]]]
Are Joins over LSM-trees Ready: Take RocksDB as an Example
Weiping Yu Fan Wang
Xuwei Zhang
Siqiang Luo
@@
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Log-structured merge (LSM) trees (O’Neil et al., 1996) based key-value stores have gained significant traction in the industry. Notable examples include RocksDB (Facebook, 2016) at Facebook, LevelDB(Google, 2016) and BigTable (Chang et al., 2008) at Google, HBase (hba, 2016) and Cassandra (Apache, 2016) at Apache, X-Engine (Huang et al., 2019) at Alibaba, WiredTiger (Wir, 2021) at MongoDB, and Dynamo (DeCandia et al., 2007) at Amazon. These LSM-based key-value stores play crucial roles in various applications such as social media (Armstrong et al., 2013; Bortnikov et al., 2018), stream processing (Cao et al., 2013; Chen et al., 2016), and file systems (Jannen et al., 2015; Shetty et al., 2013).
@@
HBase (hba, 2016), LevelDB (Google, 2016), RocksDB (Facebook, 2016) $\times$ $\times$ $\times$ $\times$
@@
$N$ Number of entries stored in an LSM-tree entries
@@ 3.1. Secondary index types analysis
Lazy Index. Lazy Index also uses posting lists to store associated primary keys within an index entry but defers their merging. Unlike Eager Index, it merely appends a new entry to the write buffer, with index entries gradually merging during compaction at a cost of $O(L\cdot T\cdot\frac{e}{B})$. As a result, entries with the same key may exist concurrently at different LSM-tree levels, requiring a modified lookup process that examines all levels to aggregate all relevant entries. This results in a lookup cost of $O(L\cdot\lceil\frac{e}{B}\rceil)$. For instance, as Figure  2 shows, retrieving entries with join attribute ${E}$ requires extracting index entries from multiple levels to compile the complete posting list $\{6,1,8\}$. Lazy Index provides better update efficiency but incurs higher lookup overhead, suiting workloads requiring rapid index updates but tolerating infrequent point lookups.
@@ 3.2. Consistency strategy analysis
INLJ-SS Secondary index Secondary index $D(R)+D(S)+D(R^{\prime})+D(S^{\prime})$ $(N_{R^{\prime}}\cdot\frac{e_{R^{\prime}}}{B})+N_{R}((1-\epsilon_{R})Z_{0}(S^{\prime})+\epsilon_{R}\cdot Z_{1}(S^{\prime}))$
@@
SJ-SS Secondary index Secondary index $D(R)+D(S)+D(R^{\prime})+D(S^{\prime})$ $(N_{S^{\prime}}\cdot\frac{e_{S^{\prime}}}{B})+(N_{R^{\prime}}\cdot\frac{e_{R^{\prime}}}{B})$
@@ 1 Lookup $n_{i}$ in $R^{\prime}$ for the corresponding primary key $p_{x}$
2 if $p_{x}$ is $NULL$ then
3       return $NULL$
@@ 4.1. Current join methods can be suboptimal
Remark: Eager Index and Validation are underestimated in conventional LSM-tree databases. Eager Index excels with frequent joins (e.g., more than twice per 10 million updates), while Validation outperforms Synchronous when entry sizes are large (e.g., over 512 bytes on the $Face$ dataset).
@@ 5.2. Future Directions
Multi-way joins are significantly more complex than binary joins that we discussed. Nevertheless, multi-way joins can be decomposed into a sequence of binary joins (Marcus and Papaemmanouil, 2018; Marcus et al., 2019; Marcus and Papaemmanouil, [n.d.]; Justen et al., 2024; Marcus et al., 2021b) thus our derived insights are still useful. For instance, we consider a join over three tables $R\Join S\Join W$, where ($R\Join S$) aligns with our provided analysis and our insights still apply. This join produces an intermediate table ($V$) with $N_{V}$ tuples, which should be stored in an LSM-tree before joining with $W$, leading to an overhead $U_{V}$ of $O(L_{V}\cdot T\cdot N_{V}\cdot\frac{e_{V}}{B})$. It’s straightforward to make $V$ indexable to boost the subsequent join during LSM-tree construction. As a result, the costs of ($V\Join W$) for INLJ, SJ, and HJ are $U_{V}+(N_{W}\cdot\frac{e_{W}}{B})+N_{W}((1-\epsilon_{W})Z_{0}(V)+\epsilon_{W}\cdot Z_{1}(V))$, $U_{V}+(N_{V}\cdot\frac{e_{V}}{B})+5(N_{W}\cdot\frac{e_{W}}{B})$, and $U_{V}+(3N_{V}\cdot\frac{e_{V}}{B})+3(N_{W}\cdot\frac{e_{W}}{B})$, respectively. This indicates that join method selection is heavily influenced by $N_{V}$. If $N_{V}$ is much larger than $N_{W}$, the costs pertaining to $W$ are negligible, indicating INLJ is preferable as it avoids extensive I/O overhead for reading the entire table $V$. If $N_{V}$ is comparable to $N_{W}$, the selection remains consistent with our previous analysis. For small $N_{V}$, INLJ is still preferable and SJ could outperform HJ after introducing a secondary index. Meanwhile, INLJ remains the best due to the desirable lookup cost achieved by the Bloom filter. These reveal enhanced attractiveness of INLJ in multi-way joins over LSM-tree. Moreover, adding a secondary index for subsequent binary joins becomes less critical, especially when $V$ is large, as an existing index is already sufficient in many scenarios.   Additionally, many existing studies enhance join performance by optimizing binary join sequences to which our work could be adapted. Specifically, our theoretical models offer cost estimations and could be incorporated into optimization frameworks like Bao (Khazaie and Pirk, 2023) to generate optimized query plans by applying the above analysis to different join orders.. The specialized schemes and index structures like multi-way hash trie join  (Khazaie and Pirk, 2023) are beyond the scope of this paper.
@@ 5.5. Supplementary analysis on Concurrency
To further investigate the effects of potential variations in implementation, we examine the impact of concurrency in joins over LSM-trees on Face, a representative real dataset with uniform data distribution. Although concurrency does not appear to impact the I/O cost presented in Table  4, systems can still benefit from the parallel execution of join tasks which deserves further evaluation. To achieve this, we divide the key ranges approximately evenly across threads, allowing each thread to scan a different range concurrently. These scans allow HJ to build hash tables in parallel, SJ to sort data of each range before merging them, and INLJ to perform concurrent point queries on the target data table. As shown in Figure 13 (c), INLJ-P shows the best thread scaling, reducing latency by over 80% at 16 threads due to minimal thread interdependence - each thread independently scans $R$R table and performs point lookups on $S$ table. SJ-P and HJ-P show less improvement as they are memory-bound, with threads waiting for memory resources. SJ-PS gains limited benefit as threading mainly affects index building compaction, which is heavily I/O-bound. Therefore, under memory constraints, INLJ methods demonstrate better concurrency potential.

### v3

# Are Joins over LSM-trees Ready: Take RocksDB as an Example

@@ Abstract.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
@@ 1. Introduction
Log-structured merge (LSM) trees (O’Neil et al., 1996) based key-value stores have gained significant traction in the industry. Notable examples include RocksDB (Facebook, 2016) at Facebook, LevelDB(Google, 2016) and BigTable (Chang et al., 2008) at Google, HBase (27) and Cassandra (Apache, 2016) at Apache, X-Engine (Huang et al., 2019) at Alibaba, WiredTiger (72) at MongoDB, and Dynamo (DeCandia et al., 2007) at Amazon. These LSM-based key-value stores play crucial roles in various applications such as social media (Armstrong et al., 2013; Bortnikov et al., 2018), stream processing (Cao et al., 2013; Chen et al., 2016), and file systems (Jannen et al., 2015; Shetty et al., 2013).
@@
HBase (27), LevelDB (Google, 2016), RocksDB (Facebook, 2016) $\times$ $\times$ $\times$ $\times$
@@
NN Number of entries stored in an LSM-tree entries
@@ 3.1. Secondary index types analysis
Lazy Index. Lazy Index also uses posting lists to store associated primary keys within an index entry but defers their merging. Unlike Eager Index, it merely appends a new entry to the write buffer, with index entries gradually merging during compaction at a cost of $O(L\cdot T\cdot\frac{e}{B})$. As a result, entries with the same key may exist concurrently at different LSM-tree levels, requiring a modified lookup process that examines all levels to aggregate all relevant entries. This results in a lookup cost of $O(L\cdot\lceil\frac{e}{B}\rceil)$. For instance, as Figure  2 shows, retrieving entries with join attribute E{E} requires extracting index entries from multiple levels to compile the complete posting list $\{6,1,8\}$. Lazy Index provides better update efficiency but incurs higher lookup overhead, suiting workloads requiring rapid index updates but tolerating infrequent point lookups.
@@ 3.2. Consistency strategy analysis
INLJ-$S$ Secondary index Secondary index $D(R)+D(S)+D(R^{\prime})+D(S^{\prime})$ $(N_{R^{\prime}}\cdot\frac{e_{R^{\prime}}}{B})+N_{R}((1-\epsilon_{R})Z_{0}(S^{\prime})+\epsilon_{R}\cdot Z_{1}(S^{\prime}))$
@@
SJ-$S$ Secondary index Secondary index $D(R)+D(S)+D(R^{\prime})+D(S^{\prime})$ $(N_{S^{\prime}}\cdot\frac{e_{S^{\prime}}}{B})+(N_{R^{\prime}}\cdot\frac{e_{R^{\prime}}}{B})$
@@ 1 Lookup $n_{i}$ in $R^{\prime}$ for the corresponding primary key $p_{x}$
2 if $p_{x}$ is NULLNULL then
3  return NULLNULL
@@ 4.1. Current join methods can be suboptimal
Remark: Eager Index and Validation are underestimated in conventional LSM-tree databases. Eager Index excels with frequent joins (e.g., more than twice per 10 million updates), while Validation outperforms Synchronous when entry sizes are large (e.g., over 512 bytes on the FaceFace dataset).
@@ 5.2. Future Directions
Multi-way joins are significantly more complex than binary joins that we discussed. Nevertheless, multi-way joins can be decomposed into a sequence of binary joins (Marcus and Papaemmanouil, 2018; Marcus et al., 2019; Marcus and Papaemmanouil, ; Justen et al., 2024; Marcus et al., 2021b) thus our derived insights are still useful. For instance, we consider a join over three tables $R\Join S\Join W$, where ($R\Join S$) aligns with our provided analysis and our insights still apply. This join produces an intermediate table ($V$) with $N_{V}$ tuples, which should be stored in an LSM-tree before joining with $W$, leading to an overhead $U_{V}$ of $O(L_{V}\cdot T\cdot N_{V}\cdot\frac{e_{V}}{B})$. It’s straightforward to make $V$ indexable to boost the subsequent join during LSM-tree construction. As a result, the costs of ($V\Join W$) for INLJ, SJ, and HJ are $U_{V}+(N_{W}\cdot\frac{e_{W}}{B})+N_{W}((1-\epsilon_{W})Z_{0}(V)+\epsilon_{W}\cdot Z_{1}(V))$, $U_{V}+(N_{V}\cdot\frac{e_{V}}{B})+5(N_{W}\cdot\frac{e_{W}}{B})$, and $U_{V}+(3N_{V}\cdot\frac{e_{V}}{B})+3(N_{W}\cdot\frac{e_{W}}{B})$, respectively. This indicates that join method selection is heavily influenced by $N_{V}$. If $N_{V}$ is much larger than $N_{W}$, the costs pertaining to $W$ are negligible, indicating INLJ is preferable as it avoids extensive I/O overhead for reading the entire table $V$. If $N_{V}$ is comparable to $N_{W}$, the selection remains consistent with our previous analysis. For small $N_{V}$, INLJ is still preferable and SJ could outperform HJ after introducing a secondary index. Meanwhile, INLJ remains the best due to the desirable lookup cost achieved by the Bloom filter. These reveal enhanced attractiveness of INLJ in multi-way joins over LSM-tree. Moreover, adding a secondary index for subsequent binary joins becomes less critical, especially when $V$ is large, as an existing index is already sufficient in many scenarios.   Additionally, many existing studies enhance join performance by optimizing binary join sequences to which our work could be adapted. Specifically, our theoretical models offer cost estimations and could be incorporated into optimization frameworks like Bao (Khazaie and Pirk, 2023) to generate optimized query plans by applying the above analysis to different join orders.. The specialized schemes and index structures like multi-way hash trie join  (Khazaie and Pirk, 2023) are beyond the scope of this paper.
@@ 5.5. Supplementary analysis on Concurrency
To further investigate the effects of potential variations in implementation, we examine the impact of concurrency in joins over LSM-trees on Face, a representative real dataset with uniform data distribution. Although concurrency does not appear to impact the I/O cost presented in Table  4, systems can still benefit from the parallel execution of join tasks which deserves further evaluation. To achieve this, we divide the key ranges approximately evenly across threads, allowing each thread to scan a different range concurrently. These scans allow HJ to build hash tables in parallel, SJ to sort data of each range before merging them, and INLJ to perform concurrent point queries on the target data table. As shown in Figure 13 (c), INLJ-P shows the best thread scaling, reducing latency by over 80% at 16 threads due to minimal thread interdependence - each thread independently scans RRR table and performs point lookups on $S$ table. SJ-P and HJ-P show less improvement as they are memory-bound, with threads waiting for memory resources. SJ-PS gains limited benefit as threading mainly affects index building compaction, which is heavily I/O-bound. Therefore, under memory constraints, INLJ methods demonstrate better concurrency potential.